# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products with applications to Kohn-Sham density functional theory

### v4

@@ A Useful Lemmas
$\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq(\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ +((\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\zeta\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)\lVert\textbf{R}^{(i)}\rVert\,$
Now, using lemma A.3 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\!\!=\!\lVert\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{B}\textbf{R}^{(i)}\rVert\,\!\!\leq\lvert\frac{\sigma_{1}}{e}\rvert\,\zeta\lVert\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$, $h_{0}=(\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=((\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\zeta\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)\lVert\textbf{R}^{(i)}\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq(\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k})\lVert\textbf{R}^{(i)}\rVert\,$ with $\eta_{k}=2f_{k}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,g_{k-1}^{1}g_{k-1}^{0}$ and $\tilde{\eta}_{k}=\lvert\frac{\sigma_{1}}{e}\rvert\,\lVert\textbf{B}\rVert\,g_{k-1}^{0}+(f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)g_{k-1}^{1}g_{k-1}^{0}$.

### v5

@@ A Useful Lemmas
$\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq(\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ +((\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\zeta\lVert\textbf{B}\rVert\,M_{k})\lVert\textbf{R}^{(i)}\rVert\,$
Now, using lemma A.2 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\!\!=\!\lVert\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{B}\textbf{R}^{(i)}\rVert\,\!\!\leq\lvert\frac{\sigma_{1}}{e}\rvert\,\zeta\lVert\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$, $h_{0}=(\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=((\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\zeta\lVert\textbf{B}\rVert\,M_{k})\lVert\textbf{R}^{(i)}\rVert\,$ we obtain $\lVert\hat{\boldsymbol{\Delta}}_{k}^{(i)}\rVert\,\leq(\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k})\lVert\textbf{R}^{(i)}\rVert\,$ with $\eta_{k}=2f_{k}\lVert\textbf{B}^{\frac{1}{2}}\rVert\,\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,g_{k-1}^{1}g_{k-1}^{0}$ and $\tilde{\eta}_{k}=\lVert\textbf{B}^{\frac{1}{2}}\rVert\,(\lvert\frac{\sigma_{1}}{e}\rvert\,\lVert\textbf{B}\rVert\,g_{k-1}^{0}+(f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\lVert\textbf{B}\rVert\,M_{k})g_{k-1}^{1}g_{k-1}^{0})$.